# Vision Beyond Boundaries: An Initial Design Space of Domain-specific Large Vision Models in Human-robot Interaction

### v1

@@ 1. Introduction
With the astonishing development of robotics and human-computer interaction (HCI), an extraordinary progress has been shown in the field of human–robot interaction (HRI) during past years, which is now benefiting the everyday life. From robots helping with household tasks to humans and robots working together in production lines, it has a significant impact on human society. The key to research in this area is understanding the nature of interactivity and social behavior between robots and humans (Dautenhahn, 2007), which aims at endowing robots with the ability to estimate the intent of humans and to complete tasks more effectively that meet the requirements of users. Drawing from existing literature (Murphy et al., 2010; Kosuge and Hirata, 2004; Sheridan, 2016), the conceptualization of HRI has undergone iterative refinement and predominantly encompasses several core aspects: the supervisory control of robots by humans, the cooperative and collaborative engagement between humans and robots in executing domain-specific tasks, and the facilitation of social objectives through human-robot interactions, etc. Among them, visual data is an imperative component which has been widely used in the design of intuitive HRI contexts, such as gestural interaction (Fan et al., 2022a), object segmentation (Kim et al., 2010), and video tracking (Menezes et al., 2003).
Over the past decade, the field of computer vision has undergone a remarkable evolution, paralleling the rapid advancements in deep learning technologies. Leveraging the satisfying performance of deep neural networks, vision models have achieved significant breakthroughs in a plethora of context-specific tasks (Wang et al., 2023). These tasks range from object detection and image segmentation to the more complex realm of scene reconstruction. In more recent years, the sphere of computer vision has further expanded, witnessing notable progress in areas such as visual tracking, video captioning, pose estimation, and innovative content generation. Particularly noteworthy is the emergence and application of generative artificial intelligence (GenAI) across various domains, where it has demonstrated impressive results. At the forefront of GenAI technology are large language models (LLMs), exemplified by OpenAI’s ChatGPT, which has garnered considerable attention for its proficiency in handling text-based data. Following the trail blazed by LLMs, large vision models (LVMs) have gradually emerged as a parallel strand of large-scale foundational models (Rawte et al., 2023). These models are increasingly gaining traction, playing a pivotal role in shaping the future of vision-based analysis and interpretation. LVMs represent a significant leap forward in the field, offering novel perspectives and capabilities in the processing and understanding of visual data, thereby setting new benchmarks in the landscape of AI and machine learning.
The emergence of LVMs has opened the potential to address long-standing challenges in HRI through their advanced capabilities in visual perception and interpretation (Wang et al., 2023). Initially developed for image-based tasks such as object recognition and scene understanding, these models are now being applied in robotics. Equipped with advanced visual processing, they are capable of interpreting complex visual scenes and making informed decisions, as well as processing multimodal input alongside visual data. The techniques applied in developing these models are essential for further progress, and may improving robustness and efficiency. Similar to the transformative impact of LLMs on text-oriented applications (Zhang et al., 2023a), LVMs are spearheading a revolution in the field of visual data processing. However, a notable distinction arises in their applicability: while LLMs trained on internet text are generally effective across a wide range of scenarios where the language used keeps identical, the imagery used by most formal companies significantly differs from the typical visuals found on social platforms. This discrepancy underscores the necessity for LVMs that are tailored to specific domains. Customized domain-specific LVMs have shown superior efficacy in comprehending the unique and nuanced visual content relevant to particular contexts, excelling in tasks such as visual classification, object detection, and segmentation which are pervasively utilized in recent HRI architectures.
The landscape of AI, and more specifically computer vision, has undergone significant transformation in recent years, largely due to the advent of vision transformers and large, general-purpose vision models. Several LVMs have paved the first step of visual data processing revolution, such as OpenAI’s CLIP (Contrastive Language-Image Pretraining) (Radford et al., 2021), Landing AI’s LandingLens 111https://landing.ai/platform/, Google’s ViT (Dosovitskiy et al., 2020), and SWIN Transformer (Liu et al., 2021). For domain-specific LVMs that are particularly focused on specialized enterprise applications, the GPT-4V 222https://openai.com/research/gpt-4v-system-card developed by OpenAI and SAM (Segment Anything Model) (Kirillov et al., 2023) by Meta have pioneered the progression of training extensive datasets of images, empowering businesses to customize them for a variety of applications within their specific domains. Beyond improved performance, domain-specific LVMs offer several advantages compared to normal LVMs and commonly-used vision models 333https://www.solulab.com/large-vision-models/#What_are_Large_Vision_Models: reduced training costs attributable to smaller databases, enhanced accuracy within specific domains, and the unprecedented scale via extensive parameters for fine-tuning.
@@
According to Landing AI 444https://landing.ai/, the LVM revolution is trailing LLM by two or three years, while domain-specific LVMs outperform others. Thus, we foresee that new advances brought by those models are being made regularly, leading to the development of more intelligent HRI systems and allowing robots to be more seamlessly integrated into human society (Figure 1). To our knowledge, this is the first work to explore the potential of LVMs in human–robot interaction, and our contributions can be summarized as follows: Developing an initial design space with empirical evaluation that helps to explore the possible usage of domain-specific LVMs in future HRI systems.
@@ 2. Related Work
Our work is established upon the intersection between the design space theory of HCI and the current status of LVMs in HRI.
@@ 2.1. Design Space with HRI
In HCI, a design space serves as a tool that highlights the diverse possibilities in crafting a specific type of artifact, thereby enriching and facilitating the design process. It also evolves iteratively, incorporating new design potentials into its existing dimensions based on insights gained from each design iteration (MacLean et al., 1991). Many research on the exploratory design spaces of various components in HRI have been accomplished during the past decades. In 2002, Dautenhahn investigated the design space of social robots that can be believed by humans, underpinned by finding the niche spaces (Dautenhahn, 2002). Woods et al. developed a design space in identifying children’s perspectives on distinguishing robots, pointing out that the two dimensions labelled ‘Behavioural Intention’ and ‘Emotional Expression’ included are the decisive factors (Woods, 2006). An original empirical framework was proposed by Walters to probe social robots’ appearance and behaviour, which suggested that humans favour the cases when the behaviour is consistent with the robot’s appearance (Walters, 2008). Kalegina et al. conducted a survey study and characterized a design space specifying the preferences of human perception regarding multifaceted robot faces based on different contexts. They indicated that the presence or absence of certain facial features can significantly impact how users perceive the robot’s capabilities and trustworthiness, thereby influencing its appropriateness for specific tasks or environments (Kalegina et al., 2018).
@@ 2.2. LVMs in HRI: status
A considerable volume of research has been successfully conducted by integrating advanced vision models into HRI systems for specialized tasks (Kang and Han, 2023; Ong et al., 2012; Sheron et al., 2021; Dehghan et al., 2019; Sakurai et al., 2007; Martínez and Vitria, 2001; Xu et al., 2020; Lee et al., 2016; Zhao et al., 2018). A prevalent task is object detection (Hsu et al., 2018; Shieh et al., 2013; Latif et al., 2023; Maiettini et al., 2021; Saputra et al., 2019), where vision models are tailored to enhance performance, enabling robots to interact more effectively with humans through improved recognition capabilities. Azagra et al. proposed a pipeline including interaction type recognition and target object detection learned by robots based on a series of vision models (Azagra et al., 2017). Moreover, Fang et al. developed a robust hand-held object detection framework by using YOLO (Redmon et al., 2016) and KCF (Henriques et al., 2014) algorithms on RGB-D video data, for better understanding human-object interaction with a humanoid robot (Fang et al., 2018). Yu et al. exploited the renowned LSTM model (Hochreiter and Schmidhuber, 1997) for video object detection, where they tested the proposition in a space robot-human scenario with gestural interactions (Yu et al., 2022). Another commonly-designed task in recent HRI research is visual segmentation that partitions images and videos into desired fragments (Chavarria et al., 2013; Ghidary et al., 2001; Fei et al., 2023). Ückermann et al. presented a real-time 3D segmentation algorithm that robustly separates objects from unstructured and cluttered environment. The evaluation was implemented in a human-interfered robot grasping mission with a Shadow Robot Hand (Ückermann et al., 2013). Fan et al. proposed a gesture segmentation (together with recognition) architecture based on deep convolutional neural networks with high accuracy, realizing it via an HRI control module with a mechanical arm crawler (Fan et al., 2022a). Similarly, Kim et al. designed a multimodal HRI framework involving online object segmentation and gesture recognition with visual features extraction, achieving effcient and interactive object learning (Kim et al., 2010). Besides, visual tracking is another task that is frequently employed in vision-related HRI research. My et al. (My and Zell, 2013) and Putro et al. (Putro and Jo, 2018) formulated real-time face tracking systems based on camera data for seamless interaction with moving robots. Distinctively, Song et al. developed a face tracking method but by using image data within a human-mobile robot interactive environment (Song and Chen, 2004). Nonetheless, other tasks such as pose estimation, content generation, visual recognition, etc are also engaged in many HRI related studies. Yet, there is a notable scarcity of research integrating well-defined LVMs into human-robot interaction contexts, with a particular absence of domain-specific LVMs. Given the proven effectiveness of LLMs in text-based information processing, we posit that domain-specific LVMs have the potential to transform vision-based contextual analysis across various industries, leveraging their specialized domain expertise.
3. Methodology
In this section, we present the formulation of the proposed design space encapsulating three main dimensions with included components in each dimension.
@@ 3.1. HRI Contexts
HRI has emerged as a critical area intersecting robotics and HCI during past decades. As robots increasingly coexist with humans in various work environments, the creation of systems that are both safe and user-friendly has become paramount. The challenge lies in the close physical interaction between humans and robots, necessitating the integration of human behaviors into the robot’s decision-making framework. As AI technology advances, research in HRI is dual-focused: not only on ensuring the safest possible physical interactions but also on fostering socially appropriate interactions that are sensitive to cultural nuances. The objective is to develop robots capable of intuitive communication with humans, leveraging speech, gestures, and facial expressions for seamless and natural exchanges. The scope of extends to multiple domains where robotic technologies are employed, encompassing industrial automation, medical assistance, and personal companionship, among other diverse applications.
Various studies in the literature have categorized different contexts of interactions between humans and robots, yielding similar outcomes yet grounded in distinct underlying principles (Schulz et al., 2018; Bütepage and Kragic, 2017; Sheridan, 2016). In our paper, we distill the interaction contexts of HRI, a critical dimension within the design space we propose, into three primary categories based on the literature examined (Figure 2):
@@ 3.2. Vision-based tasks
The second dimension identified in our design space pertains to vision-based tasks, developed by synthesizing insights from published works on visual engagement in HRI and current categorizations of computer vision tasks (Voulodimos et al., 2018). As illustrated in Figure 2, this dimension comprises nine distinct tasks:
@@ 3.3. Specific domains
The primary objective of our paper is to highlight the significant potential of domain-specific LVMs in enhancing the efficiency of HRI systems. In comparison to generic LVMs, these domain-specific models require only about 10% to 30% of the amount of labeled data, while producing considerably fewer errors. This represents a substantial advancement in reducing computational demands and achieving improved performance. Consequently, the specific domains of LVMs form another critical dimension in our design space (as illustrated in Figure 2). We have pinpointed eight domains that are expected to encompass most application areas of HRI, both current and prospective, based on insights from published works and research trends:
@@
4. User Validation
Exploring the design space’s dimensions offers valuable insights to HRI researchers and practitioners about the diversity within the domains of LVMs. However, this exploration does not necessarily reveal how the possible variations may trigger different responses from individuals. To bridge this gap, we carried out an expert evaluation aimed at gathering empirical evidence to guide the development of HRI systems that can evoke specific reactions to LVMs. We engaged a total of 15 participants aged between 24 and 42 (7 male, 6 female, 2 prefer not to say, M = 30.33, SD = 4.66) through email invitations and personal outreach. Among these participants, we have identified their prior experience in computer vision, robotics, HRI, and the experience in proximity with HRI and computer vision models (refer to Figure 3). We found that the majority of participants have several years of relevant experience, which strengthens this empirical validation with enhanced reliability and robustness. We presented them of the design space with a brief background introduction and formulated a questionnaire to delve into the experts’ perspectives on the three individual dimensions and the design space itself.
4.1. Questionnaire Design
This study employed a questionnaire-based evaluation alongside quantitative statistical methods to assess the utility of the design space by measuring several perceived metrics motivated from existing literature (Gries, 2004; Kalegina et al., 2018; Baartmans et al., 2020; Haggerty et al., 2011). In addition to collecting basic demographic information, the questionnaire introduced to participants included an overview of the design space (as shown in Figure 2) and posed a series of targeted questions. These questions aimed to gauge participants’ perceptions on several key aspects: perceived likeability, trustworthiness, usefulness, intent to use, comprehensiveness, and system usability of the dimensions incorporated in the design space, utilizing a 7-point Likert scale. Following the questionnaire, participants underwent a system usability scale (SUS) assessment eventually, which yields six metrics in total. The questions posed were as follows:
@@ 4.2. Data Analysis
On average, participants completed each trial within approximately 15 minutes, which included a brief interview session at the end, where they were request to share their overall impressions of the design space. Each participant successfully finished the study and received a small gift as appreciation. Prior to conducting statistical analysis, we carried out normality tests and confirmed that all the metrics adhered to a normal distribution () except the SUS score. One-way ANOVA with repeated measurements () for the five metrics and the Friedman test for the SUS score were implemented to determine the statistical significance of the results, which are comprehensively depicted in following subsections and Figure 4. Overall, each dimension and the design space itself garnered satisfactory ratings across all metrics and the SUS score. Notably, the HRI context dimension achieved the highest ratings in every metric, whereas the vision-based tasks dimension registered the lowest.
4.2.1. Likeability.
@@
Participants showed an overall preference favoring the HRI contexts dimension as the most likable, closely followed by the design space itself and then the specific domains dimension. However, the differences in likeability ratings did not reach statistical significance.
4.2.2. Trustworthiness.
@@
In terms of trustworthiness, perceptions varied slightly from likeability. Here, the specific domains dimension was considered second only to the HRI contexts dimension, though the gap was marginal. Again, these differences were not statistically significant.
4.2.3. Usefulness.
@@
The vision-based tasks dimension was seen as the least motivational for future research, distinctly lagging behind the other dimensions. This assessment was supported by a significant statistical difference, as indicated by a repeated measures ANOVA (), highlighting the variance in perceived usefulness across dimensions.
4.2.4. Intent to Use.
@@
The variance in the intention for potential adoption was the smallest across all dimensions with the design space itself when compared to other five evaluated metrics. The vision-based tasks dimension was rated only slightly falling behind other dimensions in this metric, although no significance was detected.
4.2.5. Comprehensiveness.
@@
Within the whole study, the HRI contexts dimension received the highest rating for comprehensiveness, nonetheless, the vision-based tasks dimension obtained the lowest rating. Statistical significance was affirmed by the repeated measures ANOVA on the comprehensiveness scales with .
4.2.6. SUS Score.
Consistent with the majority of metrics, the HRI contexts dimension was rated obviously the highest in terms of the SUS score, followed by the design space itself, and then the specific domains dimension. Significant differences were revealed in usability scores, highlighting the distinct user experiences across these dimensions, .
4.2.7. Summary of Qualitative Feedback.

### v2

@@ 1. Introduction
The rapid advancement of robotics and human-computer interaction (HCI) has significantly progressed human–robot interaction (HRI), impacting daily life. Key research focuses on understanding interactivity and social behavior between robots and humans (Dautenhahn, 2007), aiming to enable robots to predict human intent and efficiently complete tasks. Literature (Murphy et al., 2010; Kosuge and Hirata, 2004; Sheridan, 2016) highlights HRI’s core aspects: human supervisory control of robots, cooperative task execution, and achieving social goals through interaction. Visual data is crucial in designing intuitive HRI contexts, including gestural interaction (Fan et al., 2022a), object segmentation (Kim et al., 2010), and video tracking (Menezes et al., 2003). Over the past decade, computer vision has evolved remarkably alongside advancements in deep learning. Deep neural networks have enabled significant breakthroughs in tasks like object detection, image segmentation, and scene reconstruction (Wang et al., 2023). More recently, computer vision has seen progress in visual tracking, video captioning, pose estimation, and innovative content generation driven by the emergence of generative artificial intelligence (GenAI), with impressive applications across various domains. Large language models (LLMs) such as ChatGPT, have excelled in text-based tasks, inspiring the development of large vision models (LVMs) which are now pivotal in advancing vision-based analysis and interpretation (Rawte et al., 2023).
The emergence of LVMs has revolutionized HRI by addressing long-standing challenges in visual perception and interpretation (Wang et al., 2023). Originally developed for image-based tasks like object recognition and scene understanding, LVMs are now applied in robotics, enabling the interpretation of complex scenes and informed decision-making with multimodal input. These models enhance robustness and efficiency, similar to how LLMs have transformed text-based applications (Zhang et al., 2023a). However, unlike LLMs, which perform well across varied scenarios, LVMs require domain-specific training due to discrepancies in visual content between formal and informal platforms. Customized LVMs excel in comprehending the unique and nuanced visual content relevant to particular contexts in specific HRI systems. The landscape of AI, particularly in computer vision, has transformed significantly due to the advent of vision transformers and LVMs. Pioneering models like OpenAI’s CLIP (Radford et al., 2021), Landing AI’s LandingLens 111https://landing.ai/platform/, Google’s ViT (Dosovitskiy et al., 2020), and the SWIN Transformer (Liu et al., 2021) have revolutionized visual data processing. Domain-specific LVMs, such as OpenAI’s GPT-4V 222https://openai.com/research/gpt-4v-system-card and Meta’s SAM (Kirillov et al., 2023), focus on specialized enterprise applications, enabling businesses to customize models for their specific needs. These domain-specific LVMs offer advantages over general models, including reduced training costs, enhanced accuracy, and unprecedented scalability through extensive parameter fine-tuning 333https://www.solulab.com/large-vision-models/#What_are_Large_Vision_Models.
@@
According to Landing AI444https://landing.ai/, the LVM revolution lags behind LLMs by two to three years, but domain-specific LVMs outperform general ones. We anticipate regular advancements from these models, leading to more intelligent HRI systems and seamless robot integration into human society. To our knowledge, this work is the first to explore LVMs’ potential in human–robot interaction, contributing an initial design space and empirical evaluation to guide the use of domain-specific LVMs in future HRI systems.
@@ 2.1. Design Space with HRI
In HCI, a design space is a tool that highlights diverse possibilities for crafting specific artifacts, enriching and facilitating the design process. It evolves iteratively, incorporating new potentials based on insights from each design iteration (MacLean et al., 1991). Numerous studies have explored design spaces in HRI. In 2002, Dautenhahn investigated the design space of social robots, identifying niche spaces for human trust (Dautenhahn, 2002). Woods et al. developed a design space to identify children’s perspectives on robots, highlighting ’Behavioral Intention’ and ’Emotional Expression’ as key factors (Woods, 2006). Walters proposed an empirical framework examining social robots’ appearance and behavior, finding that consistent behavior and appearance are preferred by humans (Walters, 2008). Kalegina et al. conducted a survey on human perception of robot faces, noting that certain facial features significantly impact perceptions of a robot’s capabilities and trustworthiness, influencing its appropriateness for specific tasks or environments (Kalegina et al., 2018).
@@ 2.2. LVMs in HRI: status
A considerable amount of research has integrated advanced vision models into HRI systems for specialized tasks (Kang and Han, 2023; Ong et al., 2012; Sheron et al., 2021; Dehghan et al., 2019; Sakurai et al., 2007; Martínez and Vitria, 2001; Xu et al., 2020; Lee et al., 2016; Zhao et al., 2018). Object detection is a prevalent task, with vision models enhancing performance and enabling better robot-human interaction (Hsu et al., 2018; Shieh et al., 2013; Latif et al., 2023; Maiettini et al., 2021; Saputra et al., 2019). Azagra et al. proposed a pipeline for interaction type recognition and target object detection (Azagra et al., 2017). Fang et al. developed a hand-held object detection framework using YOLO (Redmon et al., 2016) and KCF (Henriques et al., 2014) on RGB-D video data (Fang et al., 2018). Yu et al. used the LSTM model (Hochreiter and Schmidhuber, 1997) for video object detection in a space robot-human scenario with gestural interactions (Yu et al., 2022). Visual segmentation, which partitions images and videos, is another common task in HRI research (Chavarria et al., 2013; Ghidary et al., 2001; Fei et al., 2023). Ückermann et al. presented a real-time 3D segmentation algorithm for robot grasping with a Shadow Robot Hand (Ückermann et al., 2013). Fan et al. proposed a gesture segmentation and recognition architecture using deep convolutional neural networks (Fan et al., 2022a). Kim et al. designed a multimodal HRI framework for online object segmentation and gesture recognition (Kim et al., 2010). Besides, Visual tracking is also frequently used in HRI research. My et al. (My and Zell, 2013) and Putro et al. (Putro and Jo, 2018) developed real-time face tracking systems for seamless interaction with moving robots, while Song et al. created a face tracking method using image data in a human-mobile robot environment (Song and Chen, 2004). Other tasks like pose estimation, content generation, and visual recognition are also common in HRI studies. However, there is a notable lack of research integrating well-defined LVMs into human-robot interaction, especially domain-specific LVMs. Given the effectiveness of LLMs in text-based information processing, we posit that domain-specific LVMs could transform vision-based contextual analysis across various industries by leveraging their specialized expertise.
3. The Design Space
In this section, we present the formulation of the proposed design space encapsulating three main dimensions.
@@ 3.1. HRI Contexts
HRI has become a crucial field at the intersection of robotics and HCI over the past decades. As robots increasingly coexist with humans in various work environments, creating safe and user-friendly systems is essential. The challenge lies in integrating human behaviors into robots’ decision-making frameworks for close physical interactions. With AI technology advancing, HRI research focuses on ensuring safe physical interactions and fostering socially appropriate interactions sensitive to cultural nuances. The goal is to develop robots capable of intuitive communication with humans using speech, gestures, and facial expressions. HRI extends to multiple domains, including industrial automation, medical assistance, and personal companionship. Various studies have categorized different contexts of interactions between humans and robots, yielding similar outcomes based on distinct underlying principles (Schulz et al., 2018; Bütepage and Kragic, 2017; Sheridan, 2016). In our paper, we propose a new and concise categorization of HRI contexts tailored for future LVM deployment (see Figure 1).
@@ 3.2. Vision-based tasks
The second dimension identified pertains to vision-based tasks, developed by synthesizing insights from published works on visual engagement in HRI and current categorizations of computer vision tasks (Voulodimos et al., 2018) (Figure 1).
@@ 3.3. Specific domains
Compared to generic LVMs, domain-specific models require only 10% to 30% of the labeled data and produce significantly fewer errors. This advancement reduces computational demands and improves performance. We identified eight domains (Figure 1) expected to encompass most current and prospective HRI application areas, based on insights from published works and research trends.
@@
4. Evaluation
4.1. Study Design
Exploring the design space’s dimensions provides valuable insights into LVM diversity for HRI researchers and practitioners. However, it doesn’t reveal how these variations affect individual responses. To address this, we conducted an expert evaluation to gather empirical evidence for developing HRI systems that evoke specific reactions to LVMs. We engaged 15 participants aged 24 to 42 (7 male, 6 female, 2 prefer not to say, M = 30.33, SD = 4.66) through email invitations and personal outreach. Participants had several years of experience in relevant areas (see Figure 2.A), enhancing the study’s reliability. After a brief introduction to the design space, we used a questionnaire to explore experts’ perspectives on its three dimensions and the design space itself. This study employed a questionnaire-based evaluation and quantitative analysis to measure five perceived metrics from existing literature (Gries, 2004; Kalegina et al., 2018; Baartmans et al., 2020; Haggerty et al., 2011). The questionnaire included demographic information, an overview of the design space (see Figure 1), and targeted questions to gauge perceptions of likeability, trustworthiness, usefulness, intent to use, and comprehensiveness, using a 7-point Likert scale. The questions are detailed in the appendix.
@@ 4.2. Data Analysis
On average, participants completed each trial in about 15 minutes, including a brief interview session to share their opinions. All participants successfully finished the study and received a small gift as appreciation. Before conducting statistical analysis, normality tests confirmed that all metrics adhered to a normal distribution (). We used one-way ANOVA with repeated measures () for the five metrics, wherein, each dimension and the design space received satisfactory ratings across all metrics (Figure 2.B). Notably, the HRI context dimension achieved the highest ratings in every metric, while the vision-based tasks dimension received the lowest.
@@
Likeability. Participants showed an overall preference favoring the HRI contexts dimension as the most likable, closely followed by the design space itself and then the specific domains dimension. However, the differences in likeability ratings did not reach statistical significance.
@@
Trustworthiness. In terms of trustworthiness, perceptions varied slightly from likeability. Here, the specific domains dimension was considered second only to the HRI contexts dimension, though the gap was marginal. Again, these differences were not statistically significant.
@@
Usefulness. The vision-based tasks dimension was seen as the least motivational for future research, distinctly lagging behind the other dimensions. This assessment was supported by a significant statistical difference, as indicated by a repeated measures ANOVA (), highlighting the variance in perceived usefulness across dimensions.
@@
Intent to Use. The variance in the intention for potential adoption was the smallest across all dimensions with the design space itself when compared to other five evaluated metrics. The vision-based tasks dimension was rated only slightly falling behind other dimensions in this metric, although no significance was detected.
@@
Comprehensiveness. Within the whole study, the HRI contexts dimension received the highest rating for comprehensiveness, nonetheless, the vision-based tasks dimension obtained the lowest rating. Statistical significance was affirmed by the repeated measures ANOVA on the comprehensiveness scales with .
4.3. Summary of Qualitative Feedback